OpenStudy (anonymous):

What value of x makes the equation true? 6.7 = –3.2 + x A. –9.9 B. –3.5 C. 3.5 D. 9.9

OpenStudy (domebotnos):

What do you think it is?

OpenStudy (anonymous):

Uhm Oh.. 3.5 or 9.9

OpenStudy (anonymous):

9.9 probaly

OpenStudy (domebotnos):

We want to isolate the variable so if we added 3.2 on each side, what is 6.7+3.2?

OpenStudy (anonymous):

9.9

OpenStudy (domebotnos):

Right. So we would have 9.9 = x The answer is D.

OpenStudy (anonymous):

Ohhh!

OpenStudy (domebotnos):

You could also plug the answers in to see which one is right.

OpenStudy (anonymous):

Can you help with the rest I only have 3 more D: and your really good help

OpenStudy (domebotnos):

Okay let me see them.

OpenStudy (anonymous):

What value of r makes the equation true? r + (–0.16) = 0.37

OpenStudy (anonymous):

A. –0.53 B. –0.21 C. 0.21 D. 0.53

OpenStudy (domebotnos):

Okay you can make the equation easier to understand by writing it as r - 0.16 = 0.37

OpenStudy (anonymous):

Ok

OpenStudy (domebotnos):

Add 0.16 to each side. We do that to cancel the -0.16 and isolate the r variable.

OpenStudy (domebotnos):

0.16+0.37 =?

OpenStudy (anonymous):

0.53

OpenStudy (domebotnos):

And so the answer is D.

OpenStudy (domebotnos):

Any more?

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

2 more

OpenStudy (anonymous):

At 6:00 a.m., the outside temperature was –14⁰. By noon. the outside temperature had risen by 8⁰. What was the outside temperature at noon? A. –22⁰ B. –6⁰ C. 6⁰ D. 22⁰

OpenStudy (domebotnos):

The equation would be -14+8

OpenStudy (domebotnos):

-14+8=?

OpenStudy (anonymous):

-7

OpenStudy (domebotnos):

Close, it was -6.

OpenStudy (anonymous):

Oh ok

OpenStudy (domebotnos):

The answer is B.

OpenStudy (anonymous):

Yesterday, Julie’s bank account balance was –$10.28. She deposited some money into her account so that her balance is now $39.89. How much did she deposit into her account? A. $50.17 B. $39.00 C. $29.72 D. $10.00

OpenStudy (domebotnos):

-10.28 + x = 39.89

OpenStudy (domebotnos):

add the a +10.28 to each side

OpenStudy (domebotnos):

39.89 + 10.28 = ?

OpenStudy (anonymous):

50.17

OpenStudy (domebotnos):

Correct, so the answer is A.

OpenStudy (domebotnos):

Those are all the questions right?

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

Thanks so much!

OpenStudy (domebotnos):

Your welcome. Good-bye.

OpenStudy (anonymous):

Err wait D:

OpenStudy (anonymous):

It says I got 2 wrong So I didn't pass now the questions are gonna change up

OpenStudy (domebotnos):

I have to go but maybe I can come back and help.

OpenStudy (anonymous):

Im sorry, I'm wasting your time and okay